# Superlatives in Context: Explicit and Implicit Domain Restrictions for Superlative Frames

### v1

@@ 1 Introduction
Superlatives are used to express a certain type of comparison in language. They work as domain-based comparisons: an expression like “the smallest fish" means that there is a fish which is smaller than all other fish in a specific set. An interpretation of the superlative comparison requires the human or machine to identify the target, e.g., the entity or event being the max or min of a set, and the comparison set (CS), e.g., the set of entities or events against which you are comparing the target.
@@
In ex. 1 the sentence with the superlative ‘largest’ does not provide enough information to properly define the CS, except that one is comparing a fish to other fish. With the help of the previous sentence, one can restrict the comparison set to the fish that were caught by Tom and his friends at the lake. In this paper we propose that recognizing the CS hinges on identifying the relevant entities or events from context, which can appear both before or after the expression. Specifically here, the catching event is crucial in restricting the comparison set.
@@
There are many different types of superlatives, either through the way they are expressed in syntax (f.ex. adverbial vs. adjectival) or through the way they express a semantic comparison. But most works only focus on single subtypes and do not cover all forms of superlatives.
@@
Superlatives are most interesting because they, on the one hand, involve deep semantic understanding and require inference of implicit elements in text, while on the other hand being highly prevalent in many downstream applications.
@@ 2.1 The Syntax of Superlatives
Superlatives can be broadly categorized into the following superficial forms: Adjectival superlatives, such as “Mia is the tallest girl"; adverbial superlatives, such as “Most commonly, psychologists use surveys"; and other forms which are not superlatives morphologically, but still lexicalize a superlative meaning, i.e. “The main reason" Scheible (2009).
@@ 2.2 The Semantics of Superlatives
In Example 2.2 the target of the comparison (i.e., the element that has the max/min of some set) is Nemo. The target is being compared to a set of other entities, the comparison set, which in this example is defined by all the fish in the aquarium. Each item in the comparison set has a property along which it is being compared, in this example the property is size. As we seek the smallest fish, the orientation of the size comparison is negative. By inference, comparatives can convey the same sense as superlatives, e.g. “Nemo is smaller than any of the other fish."
Towards a semantic account of superlatives, Scheible (2008) defined 3 types of superlative comparisons. The Property Set Comparison is the most known superlative type, where members of the CS are being compared with respect to the property. Ex. (2) illustrates a property set comparison, where all fish in the aquarium (CS) are being compared by the size property. The Relative Set Comparison involves two interdependent set comparisons. For example, in “Of all the band members, Bob played the longest solo.", the set of ‘solos’ is being compared in terms of length. But this set is further restricted by a second set, the ‘band members playing (solos)’. The Subject-based Set Comparison is peculiar in that the CS does not consist of different entities, but instead compares the target at different states: “Bob is hungriest at noon." Here the comparison set involves Bob’s level of hungriness at different times of the day.
@@ 2.3 Implicit Elements of Superlatives
In Example 2.3a. the target ’Nemo’ is explicit, while in 2.3b. the superlative NP stands for the implicit target.
@@
In 2.3b. the head of the superlative NP ("man") is empty Elazar and Goldberg (2019) and the domain "In Europe" is implicit, while in 2.3a. constructing the CS consists of two, syntactically non-adjacent spans, i.e., men, and in Europe.
@@ 2.4 Context Restrictions
Assuming that Bob brought cupcakes to a birthday party and utters the sentence in 2.4, it is clear that not all cupcakes in the whole universe are being referred to and quite probably also not all cupcakes at the party. Rather, the expression is equivalent to “Every cupcake Bob brought to the party is gluten free.", which can be inferred from context. Besides the case of quantifiers, context dependence also occurs with pronouns or demonstratives, where the referent can only be understood through additional context (f.ex. by seeing what someone is pointing to). Ambiguous propositions (either lexically or structurally) can also be disambiguated with context Stanley and Gendler Szabó (2000). But also non-ambiguous sentences could have a positive or negative truth value depending on context. And lastly, contexts also have pragmatic functions, for example by marking irony.
@@ 3 Superlative Frames
Figure 1: An annotation example showing on the left the superlative (most), the sentence it appears in, and its previous context (shortened). On the right it shows the annotations slots (Target, DOI, CS etc.) and how they are filled given the text. Highlighted in yellow are the implicit discourse restrictions.
@@
The frames are built with a focus on annotating the semantics of the comparison and making discourse restrictions explicit. Additionally, we center the frames around events/predicates (when available), using Neo-Davidsonian semantics. This is motivated by the fact that events can also function as set restrictions for superlatives, such as for relative set comparisons. An example annotation using our scheme can be seen in Fig. 1.
@@
In the comparison set we define the set of entities or events that take part in the comparison: $CS=\{e_{1},...,e_{n}\}$. Comparison involving an event are formulated using a neo-davidsonian expression. The argument slots are labeled using VerbNet roles Schuler (2005) and filled with tokens from context. The CS in Fig. 1 consists of a pay event, with four semantic arguments: agent, asset, location and time.
@@
Sometimes superlative targets do not denote the entity at the min/max position, but instead they denote an entity at the n-th position. For example: “the second biggest Bulgarian port". In these cases we note the given rank (default is 1).
@@ 4.1 Data
In order to extract sentences containing superlatives from our chosen corpora, we POS-tag them using Stanza Qi et al. (2020) and extract all sentences containing at least one word tagged with either JJS (adjectival) or RBS (adverbial). We run the preselection by POS-tag approach on a test set from Scheible (2008), receiving a recall of 98.8%. We can therefore be sure that we are capturing most, if not all, superlatives present in a text. In terms of precision, we note that not all words tagged with JJS signal a superlative, such as, ’at least’ or ’at most’, which are proportional quantifiers. We increase precision to 99% through manual post-processing, by having annotators mark such instances as non-superlative readings.
@@ 4.2 Annotators
We hired two annotators for the task, which were provided with guidelines and training sessions. For quality assurance the authors met with the annotators in weekly meetings and discussed a subset of the annotations. Additionally, we periodically calculated IAA between the annotators and an expert (i.e. author of the paper, from here on called annotator C). The annotators were paid above minimum wage for the region. One of the annotators was a Master student in linguistic, hereafter called annotator A, and the other, annotator B, a Bachelor student in Computer Science.
@@ 4.3 Coverage
The superlative frames, described in Sec.. 3, can be used to annotate all three semantic types defined by Scheible (2008). The example in Fig. 1 annotates a relative set comparison. A property set comparison distinguishes itself from the relative one by not having (eventive) restrictions in the target and CS. The subject-based set comparison can be identified through our annotations by the use of light verbs as the event predicate. For “Bob is hungriest at noon", the target, for example, would look as follows: BE_HUNGRY(e, THEME=Bob, TIME=at noon).
@@
Figure. 2 shows how the types are distributed across the five domains. Except for the literature domain, relative superlative comparisons are the most frequent. In the literature domain on the other hand, property set comparisons occur most often. The least frequent type is the subject-based comparison, for all domains.
@@ 4.4 Dataset Analysis
Here we present an analysis of the SuperSem dataset and report statistics. Table 1 shows the general dataset counts, split by domain. The final dataset consists of more than 3k annotated superlatives and more than 1k non-superlatives, which are pos-tagged as JJS, but do not express a superlative reading. About 42% of annotated superlatives contain implicit elements and about 35% contain an event.
@@
In Figure 3 we visualize the distribution of the most frequent roles in our target and CS annotations. Agent, location and theme are the most frequently annotated VerbNet roles. We additionally allowed the use of ‘of’ as a slot designating restricting bridging relations (such as “writers OF=the ancient world").
@@ 4.4.1 IAA
To ensure annotation consistency and quality, we performed three rounds of IAA checks. For the first two checks, a set of instances was given to two annotators: 50 Amazon review instances were shown to annotator A and C, and 81 Wikipedia instances to annotators B and C. We compare the agreement on different aspects of the annotation, as seen in Tab. 2. For all categorical values we also report Cohen’s Kappa scores, which are moderate to high for the event vs. no-event and the orientation frames and fair for the implicit frame. It is worth noting that these first IAA checks were performed while annotations were on-going and that there was an annotation consolidation after the checks. This means that the scores could be seen as a lower bound for annotation agreement, which then improved after the consolidations. Higher exact match agreement was harder to achieve for some non-categorical categories, like the CS. This is mainly due to the order in which arguments are listed and the differences in argument spans and lower agreement for these categories does not necessarily indicate wrong annotation.
@@
Table 2: Inter-Annotator Agreement on Wikipedia, Reviews and a final set randomly sampled from SuperSem. event vs. none: Accuracy of choosing an event. exact …: exact match accuracy for each of these slots. event predicate: acc. of choosing the same predicate. CS text: exact match of CS text if no event was chosen. role argument: Accuracy of having an intersection over union >= 0.5 of the role argument text. Cohen’s Kappa scores are added in brackets for all categorical values.
@@ 4.4.2 Bridging and Discourse Restrictions
Discourse restrictions and bridging relations are closely related. Bridging relations are anaphoric part-of relations between two entities in a discourse Gundel et al. (1993); Hou et al. (2013). They form a subset of the different types of discourse restrictions affecting the CS of superlatives. To analyze the influence of noun phrase (NP) relations as discourse restrictors for superlatives, we annotated a subset of TNE Elazar et al. (2022). TNE annotations follow a broader definition of bridging, by connecting NPs of any relation type.
@@
Using automatic string matching we found that about 26% of the implicit superlatives in our TNE subset are restricted by noun phrase relations also present in TNE. Due to the automatic way of extracting these statistics we assume that this is a lower bound and conclude that CS are often also restricted by NP relations from discourse.
@@ 5 Computational Modeling of Superlative Semantics
As follows we want to examine the computational modeling of superlative semantics. We are interested in multiple aspects. First, we want to establish sequence-to-sequence baselines for predicting our superlative frames, when trained on SuperSem. Additionally, we want to better understand the role of context when predicting the superlative frames, and when superlatives are ambiguous. To answer those questions, we carry out 5 types of experiments: In the first experiment we learn to predict the semantic interpretation of superlatives using seq2seq models. In the second experiment we predict discourse context given the formal semantic interpretation, in order to see whether the model is able to appropriately identify context restrictions. In the third experiment we test GPT-4’s capabilities of identifying the comparison set. As a forth experiment, we test our model on a hand-crafted challenge set of ambiguous superlative expressions. And lastly, we quantify the ambiguity of superlative expressions and how context can restrict the set of possible interpretations.
@@ 5.1 Predicting the Interpretation
    single: fine-tune a model for each frame individually.
@@ Results
Table 2 displays the results. For most of the frames, the setting which includes both the superlative sentence and the additional discourse context works best. This indicates that the context contains further information needed to make the appropriate inferences for predicting superlative interpretations. We also see that, except for the property slot, training a specialized model for each slot works better than training a general model that predicts the full annotation at once. The results on eventive superlatives only (in grey), show that they form an especially challenging subset for models. The best models still do not achieve the same EM scores as the human IAA scores (Sec. 4.4.1). While these IAA scores come from a different, and smaller, set than the test set, they can still provide a reference point.
@@
Table 3: Results showing exact match accuracy (EM), intersection over union (IOU)222 and Rouge-n (R, n=1). For each semantic slot we show the performance when only trained on that specific slot and the full performance. For target and CS, highlighted in grey, we also show performance of the single model (event) tested only on the events subset. The last column shows human EM IAA scores from the last IAA round.
@@
Here, the target is implicit. The model is nonetheless able to identify which parts of the superlative interpretation are not mentioned in the given sentence and then predicts appropriate context containing this implicit information. In this case, it predicted context mentioning eagles. In 83% of the implicit test cases, the model correctly included the missing restrictions when predicting context beyond the sentence boundaries.
@@ 5.3 Superlatives and GPT-4
We test GPT-4222Accessed on: 10.01.2023 and 05.16.2024.’s ability to zero-shot interpret superlative comparisons, in natural language, and to few-shot interpret superlatives’ CS.
@@
We perform two different experiments on the test split of SuperSem. First, a zero-shot experiment where we input a single sentence containing a superlative expression and ask GPT-4 to answer the question “What is being compared to what here with the superlative?". The model’s answer is expected to be a natural language explanation of the comparison. We also experiment with adding the full context and with explicitly marking the superlative in the prompt. Second, we evaluate GPT4 in a few-shot manner, where the task is to predict the superlative frame of the CS. Specifically, we add three demonstrations to the prompt, with each demonstration capturing a different type of CS interpretation.
@@
In the first setup, our aim was to see whether the LLM is able to recognize the comparison relation triggered by the superlative. For 16% of the test set sentences GPT-4 either outputs that “There is no direct comparison being made in this sentence.", or, in rarer cases, mentions other types of comparisons present in the sentence, such as discourse relations. Generally, though, it is able to correctly recognize that the presence of a superlative expression indicates a comparison.
@@
In this excerpt (shortened for space considerations), the target is ‘Mona’ and the CS is ‘all of the characters, excluding the main ones’, which is further restricted by a love event and by the fact that these characters are from the book ‘The Four Horsemen’. GPT-4 writes: “All other characters are being compared to Mona with the superlative." While this output correctly identifies that there is a comparison between Mona and other characters, it incorrectly writes ‘all other characters’. The correct response would have excluded the main characters from the comparison. It further misses to specify discourse restrictions, such as the book title these characters appear in.
@@
Also the manual evaluation reveals lower accuracy for the few-shot prompting setup (Tab. 4). Looking at the outputs, the model sometimes seems to be able to capture the format of the frames, also of eventive CSs. Interestingly, most of the errors involve the model not being able to incorporate the relevant elements from context, such as missing a LOCATION or TIME restriction.

### v2

# Superlatives in Context: Modeling the Implicit Semantics of Superlatives

@@ Abstract
Superlatives in Context:
Modeling the Implicit Semantics of Superlatives
Valentina Pyatkinthanks: ThisworkwascompletedinpartialfulfillmentforthePhDdegreeofValentinaPyatkin.   1,2   Bonnie Webber3   Ido Dagan1   Reut Tsarfaty1  1Bar Ilan University 2Allen Institute for AI 3The University of Edinburgh valentinap@allenai.org
@@ 1 Introduction
Superlatives are used to express a certain type of comparison in language. They work as domain-based comparisons: an expression like “the smallest fish” means that there is a fish which is smaller than all other fish in a specific set. An interpretation of the superlative comparison requires the human or machine to identify the target, e.g., the entity or event being the max or min of a set, and the comparison set (CS), e.g., the set of entities or events against which you are comparing the target.
@@
In ex. 1 the sentence with the superlative ‘largest’ does not provide enough information to properly define the CS, except that one is comparing a fish to other fish. With the help of the previous sentence, one can restrict the comparison set to the fish that were caught by Tom and his friends at the lake. In this paper we propose that recognizing the CS hinges on identifying the relevant entities or events from context, which can appear both before or after the expression. Specifically here, the catching event is crucial in restricting the CS.
@@
There are many different types of superlatives, either through the way they are expressed in syntax (e.g. adverbial vs. adjectival) or through the way they express a semantic comparison. But most works only focus on single subtypes and do not cover all forms of superlatives.
@@
Superlatives are most interesting because they, on the one hand, involve deep semantic understanding and require inference of implicit elements in text, while on the other hand are highly prevalent in many downstream applications.
@@ 2.1 The Syntax of Superlatives
Superlatives can be broadly categorized into the following superficial forms: adjectival superlatives, such as “Mia is the tallest girl”; adverbial superlatives, such as “Most commonly, psychologists use surveys”; and other forms which are not superlatives morphologically, but still lexicalize a superlative meaning, i.e. “The main reason” Scheible (2009).
@@ 2.2 The Semantics of Superlatives
In Example 2.2 the target of the comparison (i.e., the element that has the max/min of some set) is Nemo. The target is being compared to a set of other entities, the comparison set, which in this example is defined by all the fish in the aquarium. Each item in the comparison set has a property along which it is being compared, in this example the property is size. As we seek the smallest fish, the orientation of the size comparison is negative. By inference, comparatives can convey the same sense as superlatives, e.g. “Nemo is smaller than any of the other fish.”
Towards a semantic account of superlatives, Scheible (2012) defined 3 types of superlative comparisons. The Property Set Comparison is the most known superlative type, where members of the CS are being compared with respect to the property. Ex. (2) illustrates a property set comparison, where all fish in the aquarium (CS) are being compared by the size property. The Relative Set Comparison involves two interdependent set comparisons. For example, in “Of all the band members, Bob played the longest solo.”, the set of ‘solos’ is being compared in terms of length. But this set is further restricted by a second set, the ‘band members playing (solos)’. The Subject-based Set Comparison is peculiar in that the CS does not consist of different entities, but instead compares the target at different states: “Bob is hungriest at noon.” Here the comparison set involves Bob’s level of hungriness at different times of the day.
@@ 2.3 Implicit Elements of Superlatives
In 2.3b. the head of the superlative NP (”man”) is empty Elazar and Goldberg (2019) and the domain ”In Europe” is implicit, while in 2.3a. constructing the CS consists of two, syntactically non-adjacent spans, i.e., men, and in Europe.
@@ 2.4 Context Restrictions
Assuming that Bob brought cupcakes to a birthday party and utters the sentence in 2.4, it is clear that not all cupcakes in the whole universe are being referred to and quite probably also not all cupcakes at the party. Rather, the expression is equivalent to “Every cupcake Bob brought to the party is gluten free.”, which can be inferred from context.
Besides the case of quantifiers, context dependence also occurs with pronouns or demonstratives, such as “This book is really good.”, where the referent can only be understood through additional context (e.g. by seeing what someone is pointing to). Ambiguous propositions (either lexically or structurally) can also be disambiguated with context Stanley and Gendler Szabó (2000). But also non-ambiguous sentences could have a positive or negative truth value depending on context. And lastly, contexts also have pragmatic functions, for example by marking irony.
@@ 3 Superlative Frames
Figure 1: An annotation example showing on the left the superlative (most), the sentence it appears in, and its previous context (shortened). On the right it shows the annotation slots (Target, DOI, CS etc.) and how they are filled given the text. Highlighted in yellow are the implicit discourse restrictions.
@@
The frames are built with a focus on annotating the semantics of the comparison and making discourse restrictions explicit. Additionally, we center the frames around events/predicates, when available, using Neo-Davidsonian semantics. This is motivated by the fact that events can also function as set restrictions for superlatives, such as for relative set comparisons. When no event is restricting the superlative, we annotate the restricting noun phrases. An example annotation using our scheme can be seen in Fig. 1.
@@
In the comparison set we define the set of entities or events that take part in the comparison: $CS=\{e_{1},...,e_{n}\}$. Comparisons involving an event are formulated using a neo-davidsonian expression. The argument slots are labeled using VerbNet roles Schuler (2005) and filled with tokens from context. The CS in Fig. 1 consists of a pay event, with four semantic arguments: agent, asset, location and time.
@@
Sometimes superlative targets do not denote the entity at the min/max position, but instead they denote an entity at the n-th position. For example: “the second biggest Bulgarian port”. In these cases we note the given rank (default is 1).
@@ 4.1 Data
In order to extract sentences containing superlatives from our chosen corpora, we POS-tag them using Stanza Qi et al. (2020) and extract all sentences containing at least one word tagged with either JJS (adjectival) or RBS (adverbial). We run the preselection by POS-tag approach on a test set from Scheible (2008), receiving a recall of 98.8%. We can therefore be sure that we are capturing most, if not all, superlatives present in a text. In terms of precision, we note that not all words tagged with JJS signal a superlative, such as, ‘at least’ or ‘at most’, which are proportional quantifiers. We increase precision to 99% through manual post-processing, by having annotators mark such instances as non-superlative readings.
@@ 4.2 Annotators
We hired two annotators for the task, who were provided with guidelines and training sessions. For quality assurance the authors met with the annotators in weekly meetings and discussed a subset of the annotations. Additionally, we periodically calculated IAA between the annotators and an expert (i.e. author of the paper, from here on called annotator C). The annotators were paid above minimum wage for the region. One of the annotators was a Master’s student in linguistics, hereafter called annotator A, and the other, annotator B, a Bachelor’s student in Computer Science.
@@ 4.3 Coverage
The superlative frames, described in Sec. 3, can be used to annotate all three semantic types defined by Scheible (2012). The example in Fig. 1 annotates a relative set comparison. A property set comparison distinguishes itself from the relative one by not having restrictions, which can either be events or noun phrases, in the target and CS. The subject-based set comparison can be identified through our annotations by the use of light verbs as the event predicate. For “Bob is hungriest at noon”, the target, for example, would look as follows: BE_HUNGRY(e, THEME=Bob, TIME=at noon).
@@
Figure 2 shows how the types are distributed across the five domains. Except for the literature domain, relative superlative comparisons are the most frequent. In the literature domain on the other hand, property set comparisons occur most often. The least frequent type is the subject-based comparison, for all domains.
@@ 4.4 Dataset Analysis
Here we present an analysis of the SuperSem dataset and report statistics. Table 1 shows the general dataset counts, split by domain. The final dataset consists of more than 3k annotated superlatives and more than 1k non-superlatives, which are pos-tagged as JJS, but do not express a superlative reading (such as ‘most’ being used as a quantifier). About 42% of annotated superlatives contain implicit elements and about 35% contain an event.
@@
In Figure 3 we visualize the distribution of the most frequent roles in our target and CS annotations. Agent, location and theme are the most frequently annotated VerbNet roles. We additionally allowed the use of ‘of’ as a slot designating restricting bridging relations (such as “writers OF=the ancient world”).
@@ 4.4.1 IAA
To ensure annotation consistency and quality, we performed three rounds of IAA checks. For the first two checks, a set of instances was given to two annotators: 50 Amazon review instances were shown to annotator A and C, and 81 Wikipedia instances to annotators B and C. We compare the agreement on different aspects of the annotation, as seen in Tab. 2. For all categorical values we also report Cohen’s Kappa scores, which are moderate to high for the event vs. no-event and the orientation frames and fair for the implicit frame. It is worth noting that these first IAA checks were performed while annotations were on-going and that there was an annotation consolidation after the checks. This means that the scores could be seen as a lower bound for annotation agreement, which then improved after the consolidations. Higher exact match agreement was harder to achieve for some non-categorical categories, like the CS. This is mainly due to the order in which arguments are listed and the differences in argument spans (i.e. determiners being included or excluded) and lower agreement for these categories does not necessarily indicate wrong annotation.
@@
Table 2: Inter-Annotator Agreement on Wikipedia, Reviews and a final set randomly sampled from SuperSem. event vs. none: Accuracy of choosing an event. exact …: exact match accuracy for each of these slots. event predicate: acc. of choosing the same predicate. CS text: exact match of CS text if no event was chosen. role argument: Accuracy of having an intersection over union (IOU) >= 0.5 of the role argument text. Cohen’s Kappa scores are added in brackets for all categorical values.
@@ 4.4.2 Bridging and Discourse Restrictions
Discourse restrictions and bridging relations are closely related. Bridging relations are anaphoric relations, such as part-of, between two entities in a discourse Gundel et al. (1993); Hou et al. (2013). They form a subset of the different types of discourse restrictions affecting the CS of superlatives. To analyze the influence of noun phrase (NP) relations as discourse restrictors for superlatives, we annotated a subset of TNE Elazar et al. (2022). TNE annotations follow a broader definition of bridging, by connecting NPs of any relation type.
@@
Using string matching we found that about 26% of the implicit superlatives in our TNE subset are restricted by noun phrase relations also present in TNE. Due to the automatic way of extracting these statistics we assume that this is a lower bound and conclude that CS are often also restricted by NP relations from discourse.
@@ 5 Computational Modeling of Superlative Semantics
In what follows we want to examine the computational modeling of superlative semantics. We are interested in multiple aspects. First, we want to establish sequence-to-sequence baselines for predicting our superlative frames, when trained on SuperSem. Additionally, we want to better understand the role of context when predicting the superlative frames, and when superlatives are ambiguous. To answer those questions, we carry out 5 types of experiments: In the first experiment we learn to predict the semantic interpretation of superlatives using seq2seq models. In the second experiment we predict discourse context given the formal semantic interpretation, in order to see whether the model is able to appropriately identify context restrictions. In the third experiment we test GPT-4’s capabilities of identifying the comparison set. As a fourth experiment, we test our model on a hand-crafted challenge set of ambiguous superlative expressions. And lastly, we quantify the ambiguity of superlative expressions and how context can restrict the set of possible interpretations.
@@ 5.1 Predicting the Interpretation
    single: fine-tune a model for each slot individually.
@@ Results
Table 3 displays the results. For most of the frames, the setting which includes both the superlative sentence and the additional discourse context works best. This indicates that the context contains further information needed to make the appropriate inferences for predicting superlative interpretations. We also see that, except for the property slot, training a specialized model for each slot works better than training a general model that predicts the full annotation at once. The results on eventive superlatives only (in grey), show that they form a specially challenging subset for models. The best models still do not achieve the same EM scores as the human IAA scores (Sec. 4.4.1). While these IAA scores come from a different, and smaller, set than the test set, they can still provide a reference point.
@@
Table 3: Results showing exact match accuracy (EM), intersection over union (IOU) and Rouge-n (R, n=1). For each semantic slot we show the performance when only trained on that specific slot and the full performance. For target and CS, highlighted in grey, we also show performance of the single model (event) tested only on the events subset. The last column shows human EM IAA scores from the last IAA round.
@@
Here, the target is implicit. The model is nonetheless able to identify which parts of the superlative interpretation are not mentioned in the given sentence and then predicts appropriate context containing this implicit information. In this case, it predicted context mentioning eagles222‘Then the eagles swooped down and snatched him up, and he flew away […]”. In 83% of the implicit test cases, the model correctly included the missing restrictions when predicting context beyond the sentence boundaries.
@@ 5.3 Superlatives and GPT-4
We test GPT-4333Accessed on: 10.01.2023 and 05.16.2024.’s ability to zero-shot interpret superlative comparisons, in natural language, and to few-shot interpret superlatives’ CS.
@@
We perform two different experiments on the test split of SuperSem. First, a zero-shot experiment where we input a single sentence containing a superlative expression and ask GPT-4 to answer the question “What is being compared to what here with the superlative?”. The model’s answer is expected to be a natural language explanation of the comparison. We also experiment with adding the full context and with explicitly marking the superlative in the prompt. Second, we evaluate GPT4 in a few-shot manner, where the task is to predict the superlative frame of the CS. Specifically, we add three demonstrations to the prompt, with each demonstration capturing a different type of CS interpretation.
@@
In the first setup, our aim was to see whether the LLM is able to recognize the comparison relation triggered by the superlative. For 16% of the test set sentences GPT-4 either outputs that “There is no direct comparison being made in this sentence.”, or, in rarer cases, mentions other types of comparisons present in the sentence, such as discourse relations. Generally, though, it is able to correctly recognize that the presence of a superlative expression indicates a comparison.
@@
In this excerpt (shortened for space considerations), the target is ‘Mona’ and the CS is ‘all of the characters, excluding the main ones’, which is further restricted by a love event and by the fact that these characters are from the book ‘The Four Horsemen’. GPT-4 writes: “All other characters are being compared to Mona with the superlative.” While this output correctly identifies that there is a comparison between Mona and other characters, it incorrectly writes ‘all other characters’. The correct response would have excluded the main characters from the comparison. It further misses to specify discourse restrictions, such as the book title these characters appear in.
@@
In addition, the manual evaluation reveals lower accuracy for the few-shot prompting setup (Tab. 4). Looking at the outputs, the model sometimes seems to be able to capture the format of the frames, also of eventive CSs. Interestingly, most of the errors involve the model not being able to incorporate the relevant elements from context, such as missing a LOCATION or TIME restriction.